OpenStudy (megannicole51):

A rod of length 3 meters with density 1+x^2 g/m is positioned along with positive x-axis, with its left end at the origin. Find the total mass and the center of mass of the rod.

OpenStudy (a_clan):

Mass = Density * Length (for a linear object) Center of mass is at the center of the entire length of the linear object.

OpenStudy (dumbcow):

\[\large mass = \int\limits_0^3 (1+x^{2}) dx\] center of mass is at point on rod where equal mass is on either side let "m" be center of mass \[\large \int\limits_0^m (1+x^{2}) dx = \frac{1}{2}mass\]